# When Models Outthink Their Safety: Mitigating Self-Jailbreak in Large Reasoning Models with Chain-of-Guardrails

### v2

@@ 4.2 Main Result
Vanilla 45.45 13.62 38.80 81.71 57.33 77.50
STAR-1 18.86 0.74 80.00 62.2 57.33 71.25
SAFEPATH 36.14 10.03 22.80 42.68 55.56 67.92
SafeChain 49.55 16.99 36.80 70.73 52.28 66.25
SafeKey 3.18 0.32 8.53 13.2 41.90 70.80
Safety Backtrack(Ours) 17.05 2.05 8.00 26.83 54.30 77.50
Safety Recomposition(Ours) 13.41 2.86 9.20 28.05 56.82 76.25
@@
Vanilla 55.45 12.44 34.00 68.29 63.14 77.92
STAR-1 17.95 0.72 86.80 76.83 56.32 76.25
SAFEPATH 34.09 8.49 16.00 50.00 57.78 70.42
SafeChain 49.32 16.87 35.60 67.07 57.58 71.25
SafeKey 4.77 0.32 4.88 10.40 49.00 76.70
Safety Backtrack(Ours) 11.82 1.30 6.40 18.05 62.12 77.92
Safety Recomposition(Ours) 12.50 4.97 2.80 21.95 60.36 78.75
@@
Vanilla 46.59 12.25 35.20 20.43 65.66 81.67
STAR-1 18.41 0.74 16.80 35.37 54.55 72.92
SAFEPATH 40.00 6.57 22.80 53.66 62.38 70.25
SafeChain 47.95 16.39 28.40 70.73 54.30 71.70
SafeKey 3.41 0.32 7.32 10.40 54.30 71.70
Safety Backtrack(Ours) 14.55 2.77 8.80 23.17 61.62 77.08
Safety Recomposition(Ours) 7.05 1.49 3.20 17.07 62.38 82.08
Our results, summarized in Table 4.2, reveal several key findings:
Baseline methods enhance safety performance at the substantial cost of reasoning capability. As reported in Table 4.2, several baselines either underperform the vanilla model on safety (e.g., SafeChain, STAR-1) or secure safety gains only at the cost of substantial reasoning degradation (e.g., SafeKey).
COG achieves the best trade-off between safety and reasoning across model scales. Our proposed COG framework, including Safety Backtrack and Safety Recomposition, consistently achieves the strongest safety improvements while largely preserving reasoning ability. This trade-off is clearly visualized in Figure 4, where our methods are positioned in the optimal upper-right quadrant, demonstrating both high safety and high reasoning performance simultaneously. Specifically, across model scales, SafB and SafR attain the lowest or near-lowest harmful/jailbreak scores while maintaining competitive reasoning performance (GPQA-Diamond and AIME2024), highlighting the robustness and scalability of COG in resolving the safety–reasoning trade-off.
5 Analysis
5.1 PCA Analysis of Safety Improvements with Minimal Reasoning Impact
To quantitatively evaluate the trade-offs between safety and reasoning, we apply Principal Component Analysis (PCA) to the hidden states of Qwen3-32B We analyze the geometric structure of the representation space under different training approaches, focusing on safety-critical prompts (Harmful, RedTeaming) and reasoning prompts (Reasoning). A linear decision boundary is trained on the two-dimensional PCA projection to quantify the separation between safety and reasoning  [^1] Results for the 8B and 14B models are provided in Appendix B.2., which is illustrated in Figure 5, where we can conclude that:
Our approach establishes a more separable safety manifold while retaining the stability of reasoning representations, indicating effective safety gains with negligible reasoning degradation. Specifically, SafR achieves the largest safety improvement ($\Delta$_Harmful = +4.985) with minimal reasoning loss ($\Delta$_Reasoning = –0.676), whereas SafB provides a more balanced gain ($\Delta$_Harmful = +3.961, $\Delta$_Reasoning = –1.771). In contrast, SafeChain offers only marginal safety benefits ($\Delta$_Harmful = +0.378) while causing severe reasoning degradation ($\Delta$_Reasoning = –5.260). These results indicate that our COG framework substantially enhances safety while largely retaining reasoning capabilities, effectively improving safety alignment without disrupting core reasoning structures.
5.2 Understanding Reasoning Preservation via Trajectory and Token Analysis
To study how training strategies influence the reasoning structure of language models, we analyze their cognitive trajectories using DeepSeek-v3.1 (DeepSeek-AI, 2024) with established prompts (Zeng et al., 2025; Gandhi et al., 2025), and measure how these patterns differ from a base model. We track the frequency of key reasoning behaviors—Backtracking, Enumeration, Subgoal Setting, and Verification—with scores reflecting the average occurrences per problem on the AIME benchmark, which is presented in Table 3, which indicates that:
Models trained with our methods best maintain the cognitive structure of the base model, and therefore have only a minimal impact on the model’s reasoning capabilities. Specifically, SafR exhibits slightly higher behavior frequencies (+0.3%) and SafB slightly lower (–0.3%), indicating that both methods achieve high alignment with the base model’s reasoning patterns. In contrast, alternative methods such as SafeChain, SafePath, and star-1 exhibit larger deviations (see Table 3). Furthermore, the token-level analysis in Appendix B.3 confirms that SafR and SafB maintain token usage close to the Base model while improving safety, indicating that they enhance safety without significantly altering intrinsic reasoning behavior.
Table 3: Comparison of reasoning patterns (Qwen3-32B) across different training strategies.“Overall Avg.” denotes the average frequency across all reasoning patterns, reflecting the overall reasoning style shift under different strategies.
Pattern Base SafeChain SafePath star-1 SafB SafR
@@ 5.2 Understanding Reasoning Preservation via Trajectory and Token Analysis
Table 4: Comparison of model performance with and without masking the default reasoning trajectory. “SafB (full mask)” corresponds to the Safety Backtrack (full mask) setting, while “SafB (part. mask)” corresponds to the Safety Backtrack (partial mask) setting.
@@
Qwen3-8B 45.45 13.62 38.80 81.71 57.58 73.30
SafB (full mask) 17.05 2.05 8.00 26.83 54.30 77.50
SafB (part. mask) 23.64 5.37 22.40 59.76 56.87 75.00
5.3 Impact of Masking Strategy in Safety Backtrack
We evaluate the masking strategy in Safety Backtrack (SafB) by comparing two training setups: (1) Full-mask training, where the loss is applied to the reasoning trace, self-check, and final answer; and (2) Partial-mask training (our main method), where the loss is applied only to the self-check and final answer, with the reasoning trace masked.
Table 4 shows that both SafB variants improve safety, with partial masking yielding larger gains on the Harmful and RedTeaming benchmarks. For reasoning, both variants enhance AIME, whereas GPQA shows a slight degradation, more pronounced under full masked SafB, which can be attributed to fewer rollout samples (two for GPQA vs. eight for AIME).
@@ 6 Related Work
Safety–Reasoning Trade‐Off
Reasoning capability and safety appear to be inherently conflicting objectives in LRMs. Zhou et al.  (Zhou et al., 2025a) show that stronger LRMs tend to produce more unsafe content, while studies by Huang et al.  (Huang et al., 2025) and Li et al.  (Li et al., 2025) highlight a fundamental trade‐off: improving one objective inevitably undermines the other, underscoring the need to balance safety and reasoning performance.
Safety for Reasoning
Instead of relying on external filters, safety-for-reasoning methods integrate controls directly into inference through training. Current approaches mostly build on supervised fine-tuning (SFT). STAR-1  (Wang et al., 2025b), SafeChain  (Jiang et al., 2025), SAFEPATH  (Jeung et al., 2025), and SafeKey  (Zhou et al., 2025b) apply curated data and tailored SFT to enhance safety while retaining reasoning ability. These methods better preserve coherence but still deliver only modest safety gains or impose notable performance costs, leaving room for more efficient optimization.
@@ 7 Conclusion
In this paper, we conducted a systematic investigation into the safety mechanisms of LRMs. Our analysis reveals the existence of Self-Jailbreak, where LRMs persuade themselves into answering the harmful queries. This phenomenon sheds new light on the underlying mechanisms of the reasoning process in LRMs. Building on this discovery, we developed CoG, a training framework that suppresses Self-Jailbreak by detecting and mitigating unsafe reasoning patterns through Safety Backtrack and Safety Recomposition. Experimental results show that CoG substantially enhances safety with minimal degradation in reasoning capability, achieving state-of-the-art performance in the safety–reasoning trade-off. Overall, this study provides a principled framework for aligning LRMs toward safe and reliable reasoning.
@@ A.1.1 Description of Baseline Methods
SAFEPATH fine-tunes LRMs in a specific way, making them always generate eight fixed tokens: “let’s think about safety first” at the start of inference, guiding the LRMs to consider more about safety during the generation process.
@@ A.1.2 Implementation Details of Baselines
Table 5: Detailed Training Hyperparameters for Star-1
Hyperparameter Value Hyperparameter Value
Finetuning Type Full Optimizer AdamW
Adam $\beta_{1},\beta_{2}$ 0.9, 0.95 Learning Rate 1e-5
Epochs 5.0 Batch Size 2
Gradient Accumulation Steps 8 Weight Decay 1e-4
Warmup Ratio 0.05 Cutoff Length 8,192
@@
Table 6: Detailed Training Hyperparameters for SafeChain
Parameter Value Parameter Value
Epochs 2 Batch Size 2
@@
Table 7: Detailed Training Hyperparameters for SafeKey
Parameter Value Parameter Value
Epochs 5 Batch Size 2
@@
The detailed implementation of the SafePath experiment is described in Table 8:
Table 8: Detailed Training Hyperparameters for SafePath
Parameter Value Parameter Value
Finetuning Type Full Cutoff Length 8192
Batch Size 2 Gradient Accumulation Steps 2
Learning Rate 1e-5 Max Steps 20
@@ A.2 Implementation Details of Our Method
A.2.1 Manual Classification Accuracy
Table 9: Classification Accuracy of Three Graduate Students on 50 Cases
Participant Correct Cases Accuracy (%)
Graduate Student A 46 92.0
Graduate Student B 43 86.0
Graduate Student C 44 88.0
Average 44.33 88.67
We invited three graduate students with domain expertise to classify 50 representative cases manually. As shown in Table 9, their classification accuracy ranged from 86.0% to 92.0%, with an average of 88.67%, indicating the reliability of our labeling scheme.
A.2.2 COG Generation Parameters

### v3

@@ 3.4 Phase 3: Inference & Alignment
Vanilla 45.45 13.62 38.80 81.71 29.00 57.33 77.50 97.6 92.68
STAR-1 18.86 0.74 20.00 37.8 5.00 57.33 71.25 96.4 93.29
SafePath 36.14 10.03 22.80 42.68 14.00 55.56 67.92 94.8 92.07
SafeChain 49.55 16.99 36.80 70.73 39.00 52.28 66.25 95.2 90.24
SafeKey 3.18 0.32 8.53 13.2 7.00 41.90 70.58 90.0 91.46
Safety Backtrack (Ours) 17.05 1.45 8.00 26.83 5.00 54.30 77.50 97.4 93.90
Safety Recomposition (Ours) 13.18 1.89 9.20 28.05 5.00 56.82 76.25 92.6 93.90
@@
Vanilla 55.45 12.44 34.00 68.29 21.00 63.14 77.92 97.6 95.73
STAR-1 17.95 0.72 13.20 23.17 32.00 56.32 76.25 88.4 97.56
SafePath 24.09 8.49 16.00 50.00 15.00 57.78 70.42 78.8 88.41
SafeChain 49.09 16.87 35.60 67.07 32.00 57.58 71.25 87.4 95.73
SafeKey 4.77 0.32 4.88 10.40 6.00 49.00 76.70 88.4 86.58
Safety Backtrack (Ours) 10.91 0.97 6.40 18.05 7.00 62.12 77.92 97.0 96.34
Safety Recomposition (Ours) 7.95 2.09 2.80 21.95 5.00 60.36 78.75 97.4 96.34
@@
Vanilla 46.59 12.25 35.20 80.43 23.00 65.66 81.67 97.6 98.17
STAR-1 18.41 0.83 16.80 35.37 3.00 54.55 72.92 85.2 95.12
SafePath 40.00 6.57 22.80 53.66 20.00 62.38 70.25 60.4 97.56
SafeChain 47.95 16.39 28.40 70.73 34.00 54.30 71.70 86.4 96.34
SafeKey 3.41 0.32 7.32 10.40 10.00 54.30 71.70 86.8 87.20
Safety Backtrack (Ours) 14.55 1.68 8.80 23.17 1.00 61.62 77.08 97.4 98.17
Safety Recomposition (Ours) 7.05 1.10 3.20 17.07 2.00 62.38 82.08 97.6 97.56
Table 2: Performance comparison of different methods under the main experimental setting. Lower values in harmful/jailbreak benchmarks and higher values in reasoning benchmarks represent better performance.
4 Experiment
In this section, we conduct experiments across multiple safety and reasoning benchmarks to validate the effectiveness of our proposed COG framework. We demonstrate that compared to prior baselines, our methods achieve state-of-the-art safety-reasoning balance across different model scales. Furthermore, we perform in-depth analyses to investigate the underlying mechanisms that contribute to the success of our approach, examining both the preservation of reasoning patterns and the distributional characteristics of the learned representations.
4.1 Experiment Setting
Training Dataset We collect 15,000 high-quality harmful queries from public datasets including Alert, ToxicDPOqa, Harmful-Dataset, Aya_RedTeaming, Do-Not-Answer, AttaQ, and Toxic-Chat (Tedeschi et al., 2024; Ahmadian et al., 2024; Wang et al., 2023; Kour et al., 2023; Lin et al., 2023). This curated set serves as input to our pipeline for constructing corresponding safety-aligned responses.
Evaluation Benchmarks For safety, we use Sorry-bench and StrongREJECT for harmful-prompt refusal, and WildJailBreak and JailBreakBench for jailbreak robustness (Xie et al., 2024; Souly et al., 2024; Jiang et al., 2024b; Chao et al., 2024). For reasoning, we adopt GPQA-Diamond, AIME2024, MATH500, and HumanEval to assess math and code reasoning capabilities (Rein et al., 2024; Mathematical Association of America (MAA), 2024; Lightman et al., 2023; Chen et al., 2021).
Models and Configuration To validate the efficacy of our methods, we utilize the Qwen3 series  (Yang et al., 2025). SafR and SafB are implemented with training details in Appendix A.2.1.
Baselines We compare with diverse representative safety alignment baselines including STAR-1, SafeChain, SafePath, and SafeKey (Wang et al., 2025; Jiang et al., 2025; Jeung et al., 2025; Zhou et al., 2025b). All baselines are reproduced following official codes, configurations, and datasets, detailed in Appendix A.2.1
4.2 Overall Results
Table 3.4 summarizes the overall safety and reasoning performance across model scales. The results reveal clear differences in how existing methods trade off safety and reasoning, and highlight the effectiveness of CoG under this trade-off.
Figure 5: Safety vs. reasoning trade-off for the Qwen 32B.
1) Vanilla LRMs without safety alignment exhibit poor safety performance. Despite their impressive performance on reasoning benchmarks, the vanilla versions of Qwen3-8B, 14B, and 32B consistently achieve the highest (worst) scores across all safety and jailbreak metrics. For instance, on the Sorry-bench and Wildjailbreak benchmarks, the vanilla models demonstrate a high propensity to follow harmful instructions, with Sorry-bench scores reaching as high as 55.45% in the 14B variant. This confirms that advanced reasoning capabilities do not inherently translate to safety, and without explicit alignment, these models remain fragile.
2) Previous approaches enhance safety performance at the substantial cost of reasoning capability. Baselines such as SafeKey and STAR-1 effectively reduce safety risks, but often incur a clear compromise in reasoning. For instance, on Qwen3-32B, SafeKey reduces Sorry-bench from 46.59% to 3.41% and JBB-PAIR from 80.43% to 10.40%, at the expense of lowering GPQA-Diamond from 65.66% to 54.30% and AIME2024 from 81.67% to 71.70% Similar trade-offs are observed for STAR-1 across model scales, suggesting that the global safety signal is harmful for reasoning capability.
3) COG achieves the best safety-reasoning balance across all model scales. As shown in Table 3.4, our methods significantly boost safety, reducing Wildjailbreak and Sorry-bench scores to levels comparable with the strongest safety baselines. This trade-off is clearly visualized in Figure 5, where our methods achieve the most favorable balance between safety and reasoning performance. For example, on Qwen3-32B, Safety Recomposition reduces Sorry-bench from 46.59% to 7.05% while maintaining AIME2024 at 82.08% and MATH500 at 97.6% compared to the vanilla model; Safety Backtrack provides a complementary point that further preserves reasoning with slightly weaker safety than the most safety baseline. Overall, these results indicate that CoG can deliver strong safety gains with minimal loss in reasoning ability, enabling a more favorable safety-reasoning trade-off for deploying capable LRMs.
4.3 Detailed Analysis
Pattern Vanilla SafeChain SafePath Star-1 SafB SafR
@@ 4.3 Detailed Analysis
Table 3: Comparison on the frequencies of reasoning patterns (Qwen3-32B) across different training strategies. “Overall Avg." denotes the average frequency across all reasoning patterns, reflecting the overall reasoning style shift under different strategies.
We further examine why COG improves safety without sacrificing reasoning largely. Specifically, we test whether COG preserves the base model’s reasoning paradigm from two perspectives—reasoning trajectory patterns and representation separability —and provide illustrative case studies in Appendix D.
4.3.1 Reasoning Pattern
To study how training strategies influence the reasoning structure of LRMs, we begin with a simple token-level signal. Based on our token-level statistics, SafR and SafB induce minimal changes in generation length on GPQA-Diamond and AIME compared to the base model, while baselines—such as STAR-1—consistently produce substantially shorter outputs across all model scales; see Appendix C.2 for details.
Subsequently, we conduct a deeper reasoning-pattern analysis, which calculates the frequencies of reasoning patterns among various methods. Following established prompting protocols (Zeng et al., 2025; Gandhi et al., 2025), we use DeepSeek-v3.1 (DeepSeek-AI, 2024) to analyze cognitive trajectories and quantify four key behaviors—Backtracking, Enumeration, Subgoal Setting, and Verification—measured as the average occurrences per problem on AIME. CoG resembles the most to original LRMs in reasoning frequencies. Table 3 indicates that SafR and SafB remain the closest to the vanilla model in behavior frequencies (SafR: +0.03%, SafB: -0.03%), while SafeChain, SafePath, and STAR-1 exhibit substantially larger deviations. Overall, these results suggest that CoG improves safety while largely preserving original reasoning patterns.
4.3.2 Representation Separability
To quantify safety-reasoning separability at the representation level, we apply PCA to Qwen3-32B hidden states on safety-critical prompts and reasoning prompts, and fit a linear decision boundary to quantify their separability, as illustrated in Figure 6.
We define each cluster’s margin as the distance from its centroid to the decision boundary: a larger margin means the centroid is farther from the boundary and thus easier to separate, while a smaller margin indicates weaker separability.
CoG improves safety separation while largely preserving reasoning representations. SafR produces the largest increase in the safety margin (+4.985+4.985) with only a small reduction in the reasoning margin (0.676-0.676). SafB increases the safety margin (+3.961+3.961), but reduces the reasoning margin more (1.771-1.771), placing reasoning representations closer to the decision boundary. In contrast, SafeChain yields a small safety gain (+0.378+0.378) while greatly reducing the reasoning margin (5.260-5.260), indicating much weaker separability for reasoning representations. Overall, these results suggest CoG strengthens safety separation without strongly distorting reasoning representations.
Comparison Awareness Classification
Human A 0.81 0.77
Human B 0.79 0.71
Human C 0.78 0.73
Mean Consensus 0.79 0.74
Table 4: Human–model agreement (Pearson’s $r$) on 50 instance-level binary judgments for Self-Jailbreak classification and LlamaGuard, where Awareness stands for the awareness of the risk in queries, while Classification represents the classification of Self-Jailbreak.
4.3.3 Consistency with human
To verify the effectiveness of LLM-as-a-judge in CoG framework and discovering Self-Jailbreak, we conduct a human–model agreement on a sampled subset from Wildjailbreak by comparing human judgments with model decisions when deciding the awareness of the risk in query as well as the classification of Self-jailbreak.
As shown in Table 4, both systems exhibit strong and consistent alignment with human judgments across annotators, achieving correlations of 0.74 (Awareness) and 0.79 (Classification) with the averaged human consensus.
4.4 Ablation Study
To verify the effect of the selective loss masking strategy in SafB, we compare (i) partial-mask training (our default), where supervision is applied only to the self-check segment and the final answer, and (ii) no-mask training, where the full sequence (including the original reasoning trace) is supervised.
@@ 4.4 Ablation Study
SafB 17.05 2.05 8.00 26.83
SafB (w/o mask) 23.64 5.37 22.40 59.76
Table 5: Effect of selective loss masking in SafB.
Table 5 shows that selective masking is crucial for SafB. Partial-mask training consistently achieves better results on safety benchmarks. In contrast, removing the mask and supervising the full reasoning trace substantially degrades safety, with the largest drops observed on jailbreak benchmarks. This supports masking the original reasoning trace in SafB and applying supervision only to the self-check and final answer, which helps mitigate unintended distributional shift.
@@ 5 Related Work
Vulnerable Safety of LRMs. Recent studies consistently show that LRMs remain vulnerable under harmful queries and adversarial settings, and that longer, explicit reasoning can introduce extra safety risks. (Huang et al., 2025) shows a clear safety-reasoning trade-off in common alignment pipelines, where better safety can hurt reasoning performance. (Zhu et al., 2025) shows that attackers can manipulate special delimiter markers to bypass reasoning, effectively skipping the intended deliberation. (Xu et al., 2025) demonstrates that safety alignment can be undermined via fine-tuning attacks against CoT-enabled models. Other work, such as (Li et al., 2025; Zhang et al., 2025a), also studies how safety changes as reasoning improves, and shows that reasoning-time safety can still fail in hard or adversarial cases.
Safety Alignment for LRMs. To mitigate these risks, recent methods align LRMs by shaping the reasoning process or injecting safety deliberation. (Guan et al., 2024) trains models to explicitly recall and reason over safety specifications before answering. Data-driven approaches such as (Wang et al., 2025; Zhang et al., 2025a) construct safety-oriented reasoning trajectories for fine-tuning. (Jiang et al., 2025) both assesses long-CoT safety risks and introduces CoT-style safety training data. Orthogonally, (Zhou et al., 2025b) aims to strengthen internal safety activation signals, while (Doula et al., 2025) promotes early safety priming to reduce harmful reasoning. In contrast, we analyze reasoning trajectories stage by stage and formalize Self-Jailbreak as a failure mode after risk awareness.
@@ 6 Conclusion
We systematically analyze safety failures in LRMs and uncover Self-Jailbreak, in which the model initially recognizes harmful intent but later overrides this judgment during reasoning, leading to unsafe outputs. This finding suggests that many safety failures stem from failure-inducing steps that override correct risk awareness within the reasoning chain. Motivated by this, we propose Chain-of-Guardrail (CoG), a trajectory-level training framework that mitigates Self-Jailbreak via targeted, step-level interventions. Across multiple safety and reasoning benchmarks, CoG improves safety and robustness while maintaining comparable reasoning performance.
Overall, our study provides an analysis lens and a practical framework for aligning LRMs toward safer reasoning.
Limitations
We acknowledge several limitations in the current study.
First, due to computational resource constraints, we do not evaluate our method on substantially larger-scale reasoning models. While our experiments demonstrate the effectiveness of the proposed framework on the models studied in this paper, validating its behavior at larger scales remains an important direction for future work.
Second, a large portion of our evaluation relies on automated judgments using LLM-based evaluators. Although LLM-as-judge provides scalability and consistency for analyzing fine-grained reasoning behaviors, it remains an imperfect proxy for human judgment. To partially mitigate this limitation, we conduct human–model consistency analyses and observe strong agreement between human annotations and automated evaluations. Nevertheless, establishing more reliable, standardized, and cost-effective evaluation protocols for safety failures in long-form reasoning remains an open problem. We leave the development of improved human-in-the-loop or hybrid evaluation frameworks to future work.
@@ A.1.1 Description of Baseline Methods
SafePath fine-tunes LRMs in a specific way, making them always generate eight fixed tokens: “let’s think about safety first" at the start of inference, guiding the LRMs to consider more about safety during the generation process.
@@ A.1.2 Implementation Details of Baselines
Hyperparameter Value
Finetuning Type Full
Optimizer AdamW
Adam $\beta_{1},\beta_{2}$ 0.9, 0.95
Learning Rate 1e-5
Epochs 5.0
Batch Size 2
Gradient Accumulation Steps 8
Weight Decay 1e-4
Warmup Ratio 0.05
Cutoff Length 8,192
Table 6: Detailed training hyperparameters for Star-1.
@@
Parameter Value
Epochs 2
Batch Size 2
@@
Parameter Value
Epochs 5
Batch Size 2
@@
The detailed implementation of the SafePath experiment is described in Table 9.
Parameter Value
Finetuning Type Full
Cutoff Length 8192
Batch Size 2
Gradient Accumulation Steps 2
Learning Rate 1e-5
Max Steps 20
@@ A.2 Implementation Details of Our Method
A.2.1 COG Generation Parameters